[Hallc_running] [New Logentry] RC report

arunts at jlab.org arunts at jlab.org
Sun Nov 14 11:15:02 EST 2021

Logentry Text:
On Friday day shift, beam was taken away for a couple of hours for pass change work for Hall A. After restoration we continued data taking. HMS Dipole ADC had read back errors which were looked into during a brief controlled access. We continued to take data over Friday swing and owl shifts. On Saturday day shift, we experienced quite a few trips (36/hr) and the beam current was lowered from 70 to 60 micro amps (https://logbooks.jlab.org/entry/3942607) which seem to have temporarily lowered the number of trips to 29/hr (https://logbooks.jlab.org/entry/3942630). It was eventually noticed by the shift crew that the energy locks were turned off in the process of delivering more stable beam. While the intention is good, it would have been nice to get a notification of the locks being turned off since stable energy is crucial for an L/T experiment and this was discussed at the 7:45 10:00 AM weekend PD meeting this morning. 

Some RF investigative work will be done on Monday and more will be discussed at the MCC meeting tomorrow. Energy measurement and beam calibrations (2:30 pm) using the arc are scheduled for Monday but would probably have to be rescheduled depending on how long the RF work will take. Other than that, we continue to take data and are on page 21/27 of the run plan.

1. HMS Dipole ADC error work: https://logbooks.jlab.org/entry/3941686, https://logbooks.jlab.org/entry/3941661
2. Spikes in beam energy: https://logbooks.jlab.org/entry/3942796


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3943200
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211114/d46edec1/attachment.html>

More information about the Hallc_running mailing list